The resistance of a wire of length 100cm diameter 0.3mm is found to be 3.0 ohms Calculate (1) the resistivity (2) the conductivity of the material of a wire


1
Expert's answer
2022-03-08T08:55:58-0500

"\rho=\frac{\pi Rd^2}{4l}=2.12\cdot 10^{-7}~Ohm\cdot m,"

"\sigma=\frac 1\rho=4.7~MSm."